OpenStudy (anonymous):

what is the apparent weight of 60kg man in a car accelerating at 5m/s^2 down a straight road. my equation W(app)=MA-MG or W(app)=M(a-g) doesnt seem to work

OpenStudy (anonymous):

It is downward movement use g-a

OpenStudy (anonymous):

I get 888,6 but the correct answer is 660

OpenStudy (anonymous):

This question kind of confuses me. Gravity influences the z-direction, while weight is a relative normal force (and varies which planent you're on and sometimes what part of the Earth you are stand at). If you're accelerating straight down, you would be moving with gravity. The Force at which you would move downwards is F=M(A-g) where g is -9.81m/s^2 therefore F=M(A+9.81

OpenStudy (anonymous):

you are not accelerating straight down but horizontally

OpenStudy (anonymous):

 If there's a ramp, then you would get the 11m/s^2 acceleration for your 660N

OpenStudy (anonymous):

the <x,-1.19> (I just made up the 1.19 to add perfectly to 11 so don't hold me on it) is the acceleration vector broken up in its x and y components. If you have an angle at which the object accelerates, you can break the acceleration down to its individual components.